# RBI-ThPhys-2025-33

### v2

@@ 3.1 The two-loop EFT
and $c_{\text{s}}^{2}=c_{\text{s},1}^{2}+c_{\text{s},2}^{2}$. The two-loop EFT Eq. (6) has counterterms $c_{\text{s},1}^{2}$ and $c_{\text{s},2}^{2}$ linked to the $\nabla^{2}\delta$ operator, $c_{\rm quad}$ to the quadratic $\nabla^{2}\delta^{2}$ term, $c_{4}$ to the quartic-derivative $\nabla^{4}\delta$ and also the stochastic $c_{\epsilon}$ parameter, totaling five free parameters. In what follows, we set $c_{\epsilon}=0$ in line with the findings of (23). Hence, we only consider the four parameters
@@ 3.3 COBRA two-loop results
We compare the two-loop power spectrum across a set of 200 randomly selected test cosmologies at $A_{s}^{*}=2.1\times 10^{-9}$ and $z^{*}=0$ within the specified parameter ranges of Table 1 for $\{\omega_{b},\omega_{c},n_{s},h\}$, so that we can explicitly assess the performance of COBRA. For these test spectra, we use 35 equally log-spaced bins in the interval $0.01,1]\,h^{*}/$Mpc. The 95th percentile of the error in each of the two-loop terms is shown in Fig. 1 for the no-wiggle and full spectra. The error is calculated relative to the full numerical result without using the COBRA basis. The horizontal dashed line indicates the $10^{-3}$ precision required by the numerical integrator, setting then a lower bound for the ratio. We show errors for different values of $N_{b}$ used in the computation of the COBRA tensors.
@@ 4.2 Likelihood and performance metrics
to consider a model fit acceptable (noting that the $1\sigma$ threshold would be $\text{FoB}<1.88$, see e.g. (59)). We also analyze the figure of merit (FoM) defined as
$\text{FoM}(k_{\text{max}})=\frac{1}{\left(\text{det}(S_{\alpha\beta})/\theta^{\alpha}_{\text{fid}}\theta^{\beta}_{\text{fid}}\right)^{1/2}}\,$, (30)
@@
with $N_{\text{dof}}$ the total number of degrees of freedom and $N_{\text{bin}}$ the number of bins included. We fix $N_{\text{dof}}=6$ for both one and two-loop analysis, corresponding to the five free cosmological parameters and the one-loop counterterm. [^6] The two-loop counterterms have no impact in the fit at low $k$, where $N_{\text{bin}}$ is small, and only start being relevant at high $k$, when $N_{\text{bin}}\gg N_{\text{dof}}$. For a more general discussion, see e.g. (60). We use the full simulation volume $V_{\text{full}}$ in the covariance when computing $\chi_{\text{red}}^{2}$, since then the covariance represents the statistical scatter in the power spectrum measurement and in that case $\chi_{\text{red}}^{2}$ should be close to unity.
@@ 4.3 Results for the comparison to simulations
Altogether, this motivates us to adopt
@@
for the two-loop EFT as baseline choices, which correspond in each case to the most conservative values obtained among the running of the counterterms, and the failure of FoB or $\chi^{2}_{\text{red}}$. Those values are indicated as vertical dashed lines in Figs. 4 and 5.
@@ Appendix C Full posteriors
Here we display the full posteriors of the one-loop and two-loop MCMC in Fig. 8. It is evident that the data is able to constrain all nuisance parameters beyond their prior ranges. We emphasize again that we do not expect the $c_{\text{s}}^{2}$ parameters to match between the one-loop and two-loop fits. We also see that at this precision, $n_{s}$ and $\omega_{b}$ cannot be constrained meaningfully, although the posterior for $n_{s}$ is markedly different from its (uniform) prior in Eq. (27). Note that for the two-loop model, the counterterms (lower right 4 by 4 triangle) are clearly correlated among each other, as are the cosmological parameters (upper left 5 by 5 triangle), but at least in the 2D marginalized posteriors the counterterms and cosmological parameters seem uncorrelated (lower left 4 by 5 rectangle).

### v3

@@ 3.1 The two-loop EFT
and $c_{\text{s}}^{2}=c_{\text{s},1}^{2}+c_{\text{s},2}^{2}$. The two-loop EFT Eq. (6) has counterterms $c_{\text{s},1}^{2}$ and $c_{\text{s},2}^{2}$ linked to the $\nabla^{2}\delta$ operator and $c_{\rm quad}$ to the quadratic $\nabla^{2}\delta^{2}$ term, i.e.
$P_{\text{quad}}=\int_{\mathbf{q}}\,F_{2}(\mathbf{q},\mathbf{k-q})P_{L}(q)P_{L}(|\mathbf{k-q}|).$ (8)
Finally, $c_{4}$ corresponds to the quartic-derivative $\nabla^{4}\delta$ and $c_{\epsilon}$ to the two-loop stochastic correction, totaling five free parameters. In what follows, we set $c_{\epsilon}=0$ in line with the findings of (23). Hence, we only consider the four parameters
@@ 3.3 COBRA two-loop results
We compare the two-loop power spectrum across a set of 200 randomly selected test cosmologies at $A_{s}^{*}=2.1\times 10^{-9}$ and $z^{*}=0$ within the specified parameter ranges of Table 1 for $\{\omega_{b},\omega_{c},n_{s},h\}$, so that we can explicitly assess the performance of COBRA. For these test spectra, we use 35 equally log-spaced bins in the interval $0.01,1]\,h^{*}/$Mpc. The 95th percentile of the error in each of the two-loop terms is shown in Fig. 1 for the no-wiggle and full spectra. The error is calculated relative to the full numerical result without using the COBRA basis. The horizontal dashed line indicates the $10^{-3}$ precision required by the CUBA numerical integrator (see the beginning of this Subsection), setting then a lower bound for the ratio. We show errors for different values of $N_{b}$ used in the computation of the COBRA tensors.
@@ 4.2 Likelihood and performance metrics
to consider a model fit acceptable (noting that the $68\%$ threshold would be $\text{FoB}<1.88$, see e.g. (59)). We also analyze the figure of merit (FoM) defined as
$\text{FoM}(k_{\text{max}})=\frac{1}{\big[\text{det}(S_{\alpha\beta}/\theta^{\alpha}_{\text{fid}}\theta^{\beta}_{\text{fid}})\big]^{1/2}}\,$, (31)
@@
with $N_{\text{dof}}$ the total number of degrees of freedom and $N_{\text{bin}}$ the number of bins included. We fix $N_{\text{dof}}=6$ for both one and two-loop analysis, corresponding to the five free cosmological parameters and the one-loop counterterm. [^6] The two-loop counterterms have no impact in the fit at low $k$, where $N_{\text{bin}}$ is small, and only start being relevant at high $k$, when $N_{\text{bin}}\gg N_{\text{dof}}$. We do find that using e.g. $N_{\text{dof}}=9$ yields somewhat low values of reduced chi-squared at low $k_{\text{max}}$, suggesting (incorrectly) that the data is being overfitted at low values of $k_{\text{max}}$. For a more general discussion, see e.g. (60). We use the full simulation volume $V_{\text{full}}$ in the covariance when computing $\chi_{\text{red}}^{2}$, since then the covariance represents the statistical scatter in the power spectrum measurement and in that case $\chi_{\text{red}}^{2}$ should be close to unity.
@@ 4.3 Results for the comparison to simulations
When determining a value of $k_{\text{max}}$, one option would be to adopt a rigorous criterion (see e.g. (59)) consisting of a combination of all three of the above criteria. However, in practice, some degree of ambiguity is always present. Hence, we simply choose a baseline scale cut that satisfies all three criteria and briefly elaborate on possible other choices at the end of this Subsection. As a baseline, we adopt
@@
for the two-loop EFT, which correspond in each case to the most conservative values obtained among the running of the counterterms, and the failure of FoB or $\chi^{2}_{\text{red}}$. Those values are indicated as vertical dashed lines in Figs. 4 and 5.
@@ Appendix C Full posteriors
Here we display the full posteriors of the one-loop and two-loop MCMC in Fig. 8. It is evident that the data is able to constrain all nuisance parameters beyond their prior ranges. We emphasize again that we do not expect the $c_{\text{s}}^{2}$ parameters to match between the one-loop and two-loop fits. We also see that at this precision, $n_{s}$ and $\omega_{b}$ cannot be constrained meaningfully, although the posterior for $n_{s}$ is markedly different from its (uniform) prior in Eq. (28). Note that for the two-loop model, the counterterms (lower right 4 by 4 triangle) are clearly correlated among each other, as are the cosmological parameters (upper left 5 by 5 triangle), but at least in the 2D marginalized posteriors the counterterms and cosmological parameters are mostly uncorrelated (lower left 4 by 5 rectangle), with the exception of e.g. $A_{s}$ and $c_{4}$ or $n_{s}$ and $c_{4}$.